Subject: APO 3.5-m Users Committee minutes 6/20/05

From: Bruce Gillespie

Submitted: Mon, 20 Jun 2005 17:58:29 -0600

Message number: 912 (previous: 911, next: 913 up: Index)

                          APO 3.5-m Users Committee Phonecon, 6/20/05

Attending:  Suzanne Hawley, Michael Strauss, Al Harper, Karl  
Glazebrook, Bruce Gillespie, Fred Hearty, John Bally, Rene Walterbos

Absent: Russet McMillan, Jon Holtzman, Don York

Minutes taken by Bruce Gillespie


User feedback, comments from Institutional representatives:

Colorado (John Bally, Fred Hearty)--John and Fred had no recent  
feedback from CU users to share (but Fred has been working with  
several NIC-FPS users from other ARC institutions lately).  John said  
that the proposed CU focal reducer for SPIcam has been on the back  
burner, but will be revived this summer.  The current plan is to  
build the system with optics behind the focal plane, and an optical  
design is being worked.  Also a medium-resolution NIR spectrograph  
add-on to NIC-FPS is in the concept phase.  The integral-field  
spectrograph, proposed last year, is on hold because of the probable  
difficulties in finding funding.  Fred reported that NIC-FPS will be  
warmed up and returned to Boulder for repair work starting on 14  
July, and return to APO around 20 August
to be ready for its engineering time on 27-28 August.  The work  
planned for it includes fixing the filter wheels, the vacuum-hold  
time, detector vignetting, and noise problems.  There is also the  
hope to add a low-res grism (salvaged from GRIMII) to the instrument.

Princeton (Michael Strauss)--Michael had nothing much to report from  
the users, other than to say that he is very happy so far with the  
NIC-FPS data he has seen.

Johns Hopkins (Karl Glazebrook)--Karl mentioned that there was a  
recent network blip experienced by a JHU remote observer, who didn't  
know how to make a long-distance call to APO to work the problem.  
That has been fixed procedurally at JHU.  BruceG mentioned that this  
loosely points to a problem at APO, in that we do not currently know  
who all the institutional systems administrators are.  We need to be  
in contact soon with these people to provide tools and instructions  
for the upcoming release and installation of TUI 1.0.

          ACTION:  All committee members should identify who their  
local systems administrator (i.e. the person who would be in charge  
of installing TUI and providing information on it to users)
is so we can contact them.  Please send the name and contact  
information to Bruce Gillespie (gillespi at

Suzanne asked if the GRB program which used JHU time to payback an  
alert recently had caused any problems and Karl said no.

NMSU (Walterbos)--Rene reported that Nancy Chanover had a very  
successful first run of the NMSU Acousto-optical Imaging Camera  
(NAIC) for narrow-band IR planetary imaging.  He also said that  
Nicole Vogt used and liked the improved user information on the APO  
web pages.  In addition, he said that a grad student, Pey-Lian Lim,  
wants to know if there is tutorial for using the echelle with TUI.  
There is user information about TUI in general, but if an instrument- 
specific tutorial is needed, it might be possible that Pey-Lian could  
help write it, since she is a frequent echelle user.  BruceG will  
talk to Jon Holtzman (Pey-Lian's advisor) about this.  Suzanne asked  
if NAIC will satisfy Nancy's needs for planetary observations, in  
lieu of certain special filters that were not possible to fit into  
NIC-FPS.  Fred said he thought that some of the filters and short-  
exposure modes in NIC-FPS do fit the bill for planetary work, but he  
agreed to talk to Nancy to be sure.

UWashington (Balick, Hawley)--BruceB reported that the department  
recently held an open meeting about APO, and no major users concerns  
were raised.  There seems to be a good level of contentment and  
satisfaction with the observatory, but not enough time was available  
in the meeting for extensive questions from users; another meeting is  

UChicago (Harper)--Al said that he had heard no recent comments from  
UChicago users, other than general discussion among the GRB folk  
about the alert program, data reductions, etc.  BruceG and Fred will  
talk to Al this Wednesday about the salvage plans for GRIMII,  
particularly on how, where, and when to remove its grism so that it  
can be used in NIC-FPS.  Suzanne mentioned that the GRB program seems  
to be running very well at the observatory level, with all pre-empted  
observers so far having been paid back with usable observing time.   
Earlier in the quarter, there was an alert about every two weeks, but  
the pace has slackened recently.  She also mentioned that the payback  
observing periods that are built into the schedule are dispensed  
optimally from the science perspective, not by institutional priority  
or sequence. Over time, this should result in the best science from  
the paybacks, and any institutional inequities should balance out.   
Also, the special schedule protection block in the proposal forms has  
been used judiciously, so far.  We will be evaluating the overall GRB  
program at the end of the quarter.


Discussion of telescope/instruments report:

Suzanne said that two highlights for the month were the continuation  
of unusually good seeing, and the implementation of the new echelle  
ICC.  BruceB asked about the possibility of using the old top end for  
mid-IR work (3 to 10 microns) as was originally envisaged for the  
site.  Al Harper said that he had no immediate interest in this (but  
"never say never"), because most of the capability for mid-IR is  
handled by the bigger telescopes at higher and drier locations.  
BruceB suggested APO still could find a niche in mid-IR monitoring  
programs, and Al said maybe.  BruceB reminded us that APO was  
originally touted as a good mid-IR site in winter.  Suzanne said this  
was a good topic for discussion by the Futures Committee, and BruceG  
mentioned that APO could be a good place for testing new mid- IR  
instruments prior to going to more remote locations.  Al said he'd  
think about it.

[Below is the monthly telescope/instrument attachment submitted to  
the committee prior to the meeting]:

      3.5-m Telescope, Instruments, and CIF Projects Highlights, 5/18  
through 6/16/05

                                  Bruce Gillespie

0) Overview

The telescope and instruments continued to give generally good  
performance during the period, with some lengthy periods of good  
observing weather and excellent seeing.  A handful of telescope and  
instrument problems were encountered and addressed, with little, if  
any, lost observing time.  One GRB alert was observed, and pay-back  
observing time to all previous GRB pre-empts was accomplished.  A new  
visiting instrument run was successfully supported.

1) Telescope

      o cleaned drive surfaces, and made minor repairs;

      o enclosure drive/servo problems (reported last month) did not  

      o some minor anomalies in NA2 guider performance addressed;

      o tested enclosure shutter emergency closure procedures;

2) Instruments

      o echelle controller upgrades implemented and tested--ready for  
user test drive

      o NIC-FPS degassed, ion gauge failure fixed;

      o NMSU Acousto-optic Imaging Camera (NAIC) visiting instrument  
used successfully on sky by
        Nancy Chanover and colleagues;

      o DIS upgrade study underway (JHU & Jon Holtzman), possible  
replacement red chip studied.

3) Engineering and CIF projects

      o New top-end:  Detailed design and fabrication cont'd, and  
tests of scattered light using new
        vane design done on telescope;

      o TUI:  Ready for release of version 1.0 shortly;

      o On-sky engineering tests:  new TUI guide s/w, fastcam data  
taken, pointing model, checked
        instrument blocks and pointing model;

      o New drive boxes:  Being made ready for installation during  
upcoming shutdowns.


Q3 schedule comments, rescheduled shutdown:

Suzanne reported that the Q3 schedule is now out.  It was somewhat  
delayed because the July shutdown needed to be shifted to a week  
later at the last minute to accommodate a people-resource issue  
between SDSS and KPNO for realuminizing the 2.5-m primary mirror.  We  
shifted our shutdown so that people critical to both our shutdown and  
the 2.5-m aluminizing were not double-booked.  Shifting our shutdown  
required some compromises but in the end was not a major problem,  
other than the extra work it took to adjust the schedule.


CIF projects status:

Suzanne said that the new top-end project, which is the big CIF  
effort this year, is moving along fairly well.  We will decide by 1  
August if the new top end will be ready for installation during  
October bright time. Also, Jon Holtzman and JHU are working up a plan  
for various DIS upgrades, which may include improved UV response, new  
gratings, and moving the dichroic split. Other DIS upgrades we are  
considering are buying a new red chip to address the fringing  
problems with the current chip, and fixing the red noise problem.   
BruceG mentioned that the new telescope drive boxes are nearly  
finished and are scheduled for installation during the summer  
shutdowns.  Also, TUI 1.0 will be released shortly, and users should  
become facile with it quickly because Remark will be retired sometime  
over the coming months.


APOLLO update:

BruceG mentioned that the APOLLO lunar laser ranging experiment is  
moving from construction to commissioning phase beginning next  
month.  In-the-enclosure and aircraft/satellite safety issues are  
being addressed, and first tests of the system are planned for July  
and August.  Once fully operational, the program will operationally  
resemble Ed Turner's lens monitoring program; two or three half-hour  
lunar ranging runs per week, typically near twilight or at the middle  
of night shift change.


CorMASS update:

BruceG reported that CorMASS is in transit back to APO from Chile. We  
plan to have this low-res NIR spectrograph available for ARC users  
during Q4 this year.  CorMASS may stay at APO into 2006. More  
information will be available for prospective 3.5-m telescope users  
over the summer.


Next phonecon will be on Monday, 19 Sept., at 8:30 AM Pacific  
Daylight Time. Agenda and other materials will be sent to the Users  
Committee members during the preceding week.  Telescope/Instrument  
monthly reports will be sent to members in July and August.  Users  
topics will also be discussed in person at the upcoming Seattle  
meeting of the Futures Committee (August 14-15, 2005).

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 912 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to